OpenStudy (anonymous):

Find sin2x, cos2x, and tan2x if cosx=3/5 and x terminates in quadrant IV.

OpenStudy (anonymous):

\[\cos x = \frac{3}{5} = \frac{adj}{hyp}\Rightarrow adj = 3, hyp = 5 \\ opp = \sqrt{hyp^2 - adj^2} = \sqrt{25 - 9} = 4 \textrm{ (also it's a 3-4-5 triangle)} \\ \sin x = \frac{opp}{hyp} = \frac{4}{5} \\ \sin 2x = 2\sin x\cos x = 2\left(\frac{4}{5}\right)\left(\frac{3}{5}\right) = \frac{24}{25} \\ \] You should be able to solve the pthers in a similar fashion.

OpenStudy (anonymous):

Thanks. But how does Sin2x become sinx cosx? I don't understand that part.

OpenStudy (whpalmer4):

Double angle formula: \[\sin 2x = 2 \sin x \cos x\]

OpenStudy (whpalmer4):

If you remember the formula for the sin of the sum of two angles: \[\sin(A+B) = \sin A \cos B + \cos A \sin B\]and let \[A = x, B=x\]\[\sin 2x = \sin(x+x) = \sin x \cos x + \cos x \sin x \]\[= \sin x \cos x + \sin x\cos x =2 \sin x\cos x\]

OpenStudy (anonymous):

Thanks for the help. This part of precalc is killin me right now.

OpenStudy (whpalmer4):

It's worth taking the time to really learn the identities...you'll make extensive use of this stuff in calculus.